what is Escape velocity?

Grade:12

7 Answers

Srikanth Neeli
11 years ago

the velocity required to escape from the the earth is called escape velocity .if you travel with 11.2km/s velocity you can escape from the earth.

 

please approve answer by clicking on yes

Revanth - -
11 years ago

DEAR SRIKANTH A

 

ESCAPE VELOCITY IS DEFINED AS THE MINIMUM VELOCITY REQUIRED BY A BODY TO ESCAPE FROM A GRAVITATIONAL FIELD THAT MAY BE OF EARTH''S OR MOON''S OR SUN''S OR EVEN A BLACK HOLE''S

please approve if u think my answer is good!!!!!!!

Swapnil Pandey
11 years ago

It is the velocity, with which an object can escape the gravitational influence of the earth. This velocity is independent of the mass of the body and its value is 11.2 km/sec approximately.

Reddy Rakesh
11 years ago

The minimum speed required by the body to escape from the earth''s gravitational field is called"Escape velocity".

upender surepally
11 years ago

Escape velocity is defined as the minimum velocty is required for the body to escape from the grvitational force of planet. The escape velocity at the Earth''s surface is about 11.2 kilometers per second (25,000 miles per hour); the escape velocity on the Moon''s surface is 2.4 kilometers per second (5,300 miles per hour).The escape velocity at the Son''s surface is about 617.3 kilometers per secon.The escape velocity at the Mercury''s surface is about 4.2 kilometers per second

G Ashwin -
11 years ago

It is the velocity required to send a body out of the gravitation field of the earth.

AAYUSH KESHAVA
11 years ago

In physics, escape velocity is the speed at which the kinetic energy plus the gravitational potential energy of an object is zero.It is the speed needed to "break free" from a gravitational field without further propulsion.

For a spherically-symmetric body, escape velocity at a certain distance is calculated by the formulav_e = \sqrt{\frac{2GM}{r}},

where G is the universal gravitational constant (G=6.67×10−11 m3 kg−1 s−2), M the mass of the planet, star or other body, and r the distance from the center of gravity.